find the length of the sides of a triangle whose vertices are (1,4),(-4,0),(3,-3)

Answer by richard1234(7193)

Use the distance formula for each pair of vertices.

E.g. d((1,4), (-4,0)) = sqrt((1-(-4))^2 + (-4-0)^2) = sqrt(25 + 16) = sqrt(41)

And so on.